OpenStudy (anonymous):

Laura has six different bracelets. She wears one on each wrist. How many different ways can she choose two bracelets to wear? 30 12 360 15

OpenStudy (anonymous):

6*5=30

OpenStudy (anonymous):

30.

OpenStudy (anonymous):

thanks both of you

OpenStudy (anonymous):

6-2=4 6*5*4/4=30

OpenStudy (anonymous):

Your Welcome!! :)

OpenStudy (anonymous):

= 6 * 5 = 30 Answer: A) 30 ways